### v1

# TrustChain: A Blockchain Framework for Auditing and Verifying Aggregators in Decentralized Federated Learning

Abstract
The server-less nature of Decentralized Federated Learning (DFL) requires allocating the aggregation role to specific participants in each federated round. Current DFL architectures ensure the trustworthiness of the aggregator node upon selection. However, most of these studies overlook the possibility that the aggregating node may turn rogue and act maliciously after being nominated. To address this problem, this paper proposes a DFL structure, called TrustChain, that scores the aggregators before selection based on their past behavior and additionally audits them after the aggregation. To do this, the statistical independence between the client updates and the aggregated model is continuously monitored using the Hilbert-Schmidt Independence Criterion (HSIC). The proposed method relies on several principles, including blockchain, anomaly detection, and concept drift analysis. The designed structure is evaluated on several federated datasets and attack scenarios with different numbers of Byzantine nodes.
Index Terms:
Federated learning, blockchain, Byzantine attacks, data poisoning, model poisoning, security.
I Introduction
The advent of Federated Learning (FL) advanced the field of distributed machine learning by introducing data decentralization as a solution to bring about data privacy and communication efficiency [1]. Despite its advantages, FL was shown to be vulnerable against a spectrum of adversaries due to its distributed nature [2, 3]. Numerous research endeavors have been dedicated to studying these threats and finding robust defense mechanisms to mitigate them.
A common perception among the majority of these studies is that the server is trustworthy, and malicious activities can potentially be initiated from the edge nodes. Thus, available defense mechanisms are mostly concerned with client activities within the FL network. Another issue of concern in FL is the fact that the server could be a single point of failure [4, 5]. From a security standpoint, this can translate into the vulnerability of the FL network when the server is breached or when the server itself is not trustworthy.
These security concerns motivated researchers to come up with decentralized architectures of FL that do not rely on a central server [6], which in turn improves the reliability and scalability of FL. In this structure, the client-side operations are similar to that of FL. However, the aggregation role is performed by one or more sets of trusted nodes. Several approaches are available for selecting the aggregator nodes in each round. While some works follow approaches as simple as following a random or specific order in selecting the aggregator nodes, others use more sophisticated technology such as Smart Contracts (SC) and proofs in blockchain to select aggregator nodes more confidently [7, 8, 9].
In the context of cyber security, several works are dedicated to securing DFL against privacy and performance attacks [10]. To mitigate privacy attacks, these structures are often combined with homomorphic encryption, differential privacy, or secure multiparty communications to secure DFL communications and minimize the risk of inference attacks [11, 12]. Dealing with performance-related attacks such as poisoning attacks, on the other hand, requires in-depth analysis of generated updates using anomaly detection and robust aggregation protocols [4, 13]. Techniques used in both of these domains are mainly inspired by previous studies in FL, as privacy-preserving and client-update evaluation are the two main security concerns in FL.
Verifying the aggregator nodes, on the other hand, is less studied in the literature. Most works in this domain assume that once a trustworthy node is selected to carry out the aggregation, it will not become rogue. This contradicts the fact that Byzantine nodes initially act honestly in the network and become dishonest once they find the opportunity to execute a malicious action. In fact, to the best of our knowledge, only a limited number of research works target this problem [14]. Available solutions mostly use a trusted committee for verifying the aggregated model [7]. Nevertheless, the trustworthiness of verifying nodes cannot be guaranteed.
To address this gap, we propose a novel approach, called TrustChain, that leverages blockchain technology to ensure the continuous trustworthiness of the aggregator node throughout the federated learning process. This method consists of two main components:
  • Pre-Selection Evaluation (PSE): When miners request downloading the blocks for aggregation, an SC is triggered to calculate a score for participating miners based on the cosine similarity of the parameters they uploaded to the blockchain and the resulting aggregated model in each round, for a window of past DFL iterations. This score indicates the tendency of each node to drift from the training path of the global model. This step ensures that only nodes with legitimate and consistent updates are considered for the aggregation role, thereby minimizing the risk of selecting a potentially malicious node.
  • Post-Aggregation Auditing (PAA): When the selected node requests to push the aggregated model to the blockchain, another SC is activated to measure the statistical independence between the aggregated model and the median of user updates using HSIC. To determine whether the obtained HSIC is safe or not, a threshold is dynamically set using the HSIC values included in previous blocks. If the new block is verified, the third SC is triggered and adds the estimated HSIC in this round to the block to facilitate the threshold estimation in the next round. This audit trail enables the detection of any malicious actions performed by the aggregator node during the aggregation process.
By integrating these two components, the proposed structure aims to enhance the security and robustness of DFL systems. The use of blockchain ensures that all transactions and updates are securely recorded, providing an additional layer of security and trust. Anomaly detection and concept drift analysis help in identifying suspicious behaviors and further strengthen the defense against malicious activities.
II Background
The literature review on the security aspects of DFL is divided into studies that analyze security threats in DFL and those designing defense mechanisms to safeguard DFL against adversaries. This section initially explains the concept of DFL and then overviews threats and defense mechanisms in this domain.
II-A Decentralized Federated Learning
DFL primarily addresses the issue of the single point of failure in FL, by removing the dependency on a reliable server to carry out the aggregation and enhance communication efficiency [15, 16]. On the client side, the process of training the local models and preparing the updates for the global aggregator is similar to that of FL. However, due to the lack of a central server in DFL, exchanging model parameters and model aggregation are performed using peer-to-peer communications or blockchain technology [10]. Integration with blockchain provides additional benefits such as traceability and immutability. Despite DFL’s security and efficiency characteristics, the design is not faultless. For example, adding blockchain into FL may leave the system vulnerable to blockchain-related security threats. More importantly, the trustworthiness of participants is more critical in DFL, as the aggregating nodes are selected among them in each round.
II-B Threats
Similar to federated learning, threats to DFL generally fall under privacy and performance-related attacks [10]. Privacy attacks primarily aim at stealing sensitive data during inference based on the communicated gradients. An example of privacy attacks in DFL is presented by [13]. On the other hand, poisoning attacks target the global model by injecting poisonous updates during the training. The attacker can use data and model poisoning to deteriorate the overall performance of the global model. For instance, [6] performs data poisoning to disrupt the global performance. This study assumes 30 percent of the nodes are malicious, and considers that this ratio remains unchanged in the experiments. The same objective can be achieved using Gradient Manipulation (GM), in which Byzantine data holders generate forged gradients sampled from a Gaussian distribution [17]. The same malicious objective is satisfied in [18] by filliping the class labels randomly on the training data. Moreover, DFL can be exposed to more subtle attacks such as backdoors as presented in [13, 19].
II-C Defenses
The majority of studies on DFL security integrate blockchain with this structure to facilitate auditing the updates and participants. Nevertheless, some studies rely solely on peer-to-peer communications and do not employ blockchain [20, 21]. These works are primarily focused on eliminating the single point of failure in conventional FL. For instance, [21] resorts to version control to perform server selection using an arbitrary pattern.
Biscotti was among the first efforts to secure DFL against privacy attacks through blockchain integration [6]. The method integrates Differential Privacy (DP) with secure aggregation as a defense against poisoning and inference attacks. Another study proposes LearningChain, a DFL structure that also uses DP and an aggregation mechanism that works based on decentralized Stochastic Gradient Descent (SGD) to eliminate Byzantine attacks [17]. One of the first efforts to demonstrate the use of smart contracts in DFL was made by [8]. This method uses smart contracts to keep track of client states and store the global model. Proof of correctness can also be used by registered users to audit the generated updates [22]. [9] uses a blockchain-based committee consensus mechanism to enhance the privacy and efficiency of DFL. The sharding approach commonly used in blockchain is used by [23] to secure the aggregation process. [7] proposed VFChain, a structure that audits trainers and generates updates using a trusted committee. The committee members are selected arbitrarily, and they use a verifying contract to verify aggregated models. Another approach proposed by [13] integrates several methods such as DP, anomaly detection, and gradient pruning to safeguard DFL against poisoning attacks. [24] tackles the issue of malicious aggregators in DFL by using zero-knowledge proofs and blockchain to enable secure and verifiable aggregation.
III Problem Formulation
Given a set of $n$ client nodes in the DFL network $C=\{c_{1},c_{2},\dots,c_{n}\}$, each node $c_{i}$ trains a local model $M_{i}$, and uploads its parameters $\theta_{i}$ to the blockchain $B$. The blockchain is formulated as $B=\{b^{1},b^{2},\dots,b^{t}\}$, where $b_{t}$ denotes the last block at round $t$ of DFL, and the others indicate the previous blocks in sequence. In this formulation, $t$ is dynamic as the size of $B$ changes in time.
In each round $t$, the last $r$ number of blocks that contain $\theta_{i}^{t}$ should be aggregated to update the global state of the model $\bar{\theta}^{t}$. Depending on the DFL structure, one node $c^{*}$ will be selected to carry out the aggregation task, which is also referred to as the selected miner of the blockchain network. Normally, $c_{a}$ aggregates $\{\theta^{t}_{i}\}_{i=1}^{r}$ into $\bar{\theta}^{t}$ and uploads it to the blockchain as $b^{t}\leftarrow\bar{\theta}^{t}$.
In this scenario, the two main attack surfaces are the client updates and the aggregated parameters. Firstly, $\theta_{i}$ uploaded to $B$ may be poisoned to satisfy a malicious goal. These threats are common with the centralized architecture of federated learning and are commonly addressed using defense mechanisms such as robust aggregation and anomaly detection. Secondly, even if the updates are filtered and the malicious updates are dropped, the node aggregating the results $c^{*}$ can still compromise the entire network by updating the global state into a malicious set of parameters.
IV Threat Models
While the spectrum of poisoning attacks is vast, this paper focuses on model poisoning attacks. We consider four poisoning attacks, namely label flipping, sign flipping, GM, and Training Objective Manipulation (TOM). When these attacks are initiated from the client side, a robust aggregator attempts to mitigate their effects. If a malicious node is selected as an aggregator, it can bypass the robust aggregator by directly feeding the poisoned parameters to the DFL. These attacks are formally explained as follows.
Label flipping
The labels of the training data are flipped to obtain corrupted parameters [25]. Let $\mathcal{D}_{i}={(x_{i},y_{i})}$ be the dataset of node $c_{i}$, where $y_{i}$ is the true label. The attacker generates a poisoned dataset $\tilde{\mathcal{D}}_{i}$ by randomly flipping labels:
$\tilde{\mathcal{D}}_{i}=\{(x_{i},\hat{y}_{i})\mid P(\hat{y}i\neq y_{i})=\gamma\}_{i=1}^{m},$ (1)
where $\gamma$ is the proportion of flipped labels, and $m$ is the number of training samples. Poisoned parameters $\tilde{\theta}_{i}$ used for attacking the DFL system are derived from this corrupted dataset:
$\tilde{\theta}_{i}\leftarrow M_{i}(\tilde{\mathcal{D}}_{i}).$ (2)
Sign flipping
The attacker reverses the sign of the gradients during parameter estimation [26]. For a gradient from node $c_{i}$, the poisoned gradient is the negative of the estimated gradients. The corresponding poisoned parameters are:
$\tilde{\theta}_{i}=\theta_{i}-\eta\cdot\text{sign}\big{(}\nabla\mathcal{L}(\theta_{i})\big{)},$ (3)
where $\eta$ is the learning rate, $\theta$ is the set of network parameters, and $\mathcal{L}$ denotes the loss function.
Gradient manipulation
During the aggregation phase, the attacker introduces noise into the aggregated parameters [27]. Let $\{\theta_{i}^{t}\}_{i=1}^{r}$ represent the set of parameters from the last $r$ blocks. The attacker modifies the aggregated update as:
$\tilde{\theta}^{t}=\frac{1}{r}\sum_{i=1}^{r}\theta_{i}^{t}+\Delta\theta,$ (4)
where $\Delta\theta$ is the noise injected by the attacker.
Training objective manipulation
The attacker modifies the training objective by adding a penalty term that deteriorates the classification performance [2]. If $\mathcal{L}(\theta)$ is the original loss function, the attacker changes it to:
~$\tilde{\mathcal{L}}(\bar{\theta}^{t})=\mathcal{L}(\bar{\theta}^{t})+\zeta\cdot\mathcal{P}(\bar{\theta}^{t}),$ (5)
where $\mathcal{P}(\bar{\theta}^{t})$ is a penalty term added by the attacker, and $\zeta$ scales the severity of the attack. We define the penalty term as:
$\mathcal{P}(\bar{\theta}^{t})=\|\nabla\mathcal{L}(\bar{\theta}^{t})-\nabla\mathcal{L}_{target}\|^{2},$ (6)
where $\mathcal{L}_{target}$ is the direction towards which the gradients are guided. Here, we choose $\mathcal{L}_{target}=-\nabla\mathcal{L}(\bar{\theta}^{t})$, which results in $\mathcal{P}(\bar{\theta}^{t})=\|2\cdot\nabla\mathcal{L}(\bar{\theta}^{t})\|^{2}$.
V Blockchain-Enabled Trustworthy Aggregation
Given that the goal of this study is to ensure the trustworthiness of the aggregator, the proposed TrustChain assumes the employed robust aggregation algorithm is reliable. Consequently, this research focuses on securing the aggregation phase by performing PSE and PAA in DFL. The process begins with the PSE algorithm evaluating the past behavior of participants using concept drift analysis to determine a set of trustworthy candidates. Then, the aggregated model is further evaluated by PAA based on the statistical independence of $\bar{\theta}^{t}$ with that of the previous block and $\theta_{i}$ received in the current round.
V-A Pre-Selection Evaluation
Input: Blockchain $B$, miner nodes $C^{*}$, number of recent blocks $q$, decay rate $\alpha$
Output: Selected aggregator node $c_{a}$
1 for $1\leq i\leq q$ do
2       Initialize score $S_{i}^{t}=0$
3 end for
4for $\forall b_{j}\in\{b_{j}\mid b_{j}\in B\}_{j=t-q+1}^{t}$ do
5       Retrieve the global state $\bar{\theta}^{j}\leftarrow b_{j}$
6       for $1\leq i\leq q$ do
7             Retrieve model parameters $\theta_{i}^{j-1}$ from $B$
8             Calculate $\operatorname{CosSim}(\theta_{i}^{j-1},\bar{\theta}^{j})$
9             Update $S_{i}^{t}\leftarrow S_{i}^{t}+\alpha^{t-j}\cdot\operatorname{CosSim}(\theta_{i}^{j},\bar{\theta}^{j+1})$
10
11       end for
12
13 end for
14Select aggregator $c_{a}\leftarrow\arg\max_{c^{*}_{i}\in C^{*}}S_{i}^{t}$
return $c_{a}$
Algorithm 1 Pre-Selection Evaluation
Figure 1: Block diagram of the proposed TrustChain structure for one DFL round at time $t$. Aggr denotes the robust aggregation algorithm. For simplicity, the diagram assumes each block number $i$ corresponds to node $c_{i}^{t}$. In this formulation, $t$ indicates the DFL iteration number. In addition, $b$ and $\bar{b}$ refer to blocks containing user updates $\theta$ and aggregated models $\bar{\theta}$, respectively.
In order to determine the trustworthiness of nodes for the aggregation phase, they are scored based on their past behavior. Specifically, we monitor the tendency of $c_{i}$ to deviate from the $\bar{\theta}$ path in terms of direction and magnitude. As detailed in Algorithm 1, this goal is achieved by measuring the cosine similarity of updates $\theta_{i}^{t-1}$ with the corresponding aggregated model for which they are used for $\bar{\theta}^{t}$.
$\operatorname{CosSim}(\theta_{i}^{t-1},\bar{\theta}^{t})=\frac{\theta_{i}^{t-1}\cdot\bar{\theta}^{t}}{\|\theta_{i}^{t-1}\|\|\bar{\theta}^{t}\|},$ (7)
where $1\leq i\leq r$. Note that we estimate the similarity for all nodes that participated in round $t-1$ regardless of the fact that not all $\theta_{i}^{t-1}$ may have been used to estimate $\bar{\theta^{t}}$. This approach results in a lower score for $c_{i}$ whose $\theta_{i}$ are disregarded or not used during the aggregation. To analyze the node behavior in a longer time span, a window of $q$ blocks is used to reveal the overall drift $c_{i}$ creates in the path of $\bar{\theta}$. Correspondingly, a score is calculated for each node at each time step.
$S_{i}^{t}=\sum_{j=t-q+1}^{t}\alpha^{t-j}\cdot\operatorname{CosSim}(\theta_{i}^{j-1},\bar{\theta}^{j}),$ (8)
where $\alpha$ is the rate of decay. The exponential decay used in this equation enables us to prioritize recent behavior and give less contribution weight to older scores. The aggregator is then selected from the pool of $k$ participating miners $C^{*}=\{c_{1}^{*},c_{2}^{*},\dots,c_{k}^{*}\}$ as:
$c_{a}=\arg\max_{c_{i}^{*}\in C^{*}}S_{i}^{t}$ (9)
V-B Post-Aggregation Auditing
Algorithm 2 details the evaluation process for validating the aggregated model. Once $c_{a}$ pushes $\bar{\theta}^{t+1}$ to the blockchain, a SC is triggered to check the statistical independence of the aggregated model with parameters used in this round $\{\theta_{i}^{t}\}_{i=1}^{r}$ and the current state of the global model $\bar{\theta}^{t}$. In this process, statistical independence is measured using HSIC. Let $A$ and $B$ be two random variables. Then, HSIC between two variables can be formally defined as:
$\begin{split}&\operatorname{HSIC}(A,B)=\\ &~{}\mathbb{E}_{ABA^{\prime}B^{\prime}}[f_{A}(A,A^{\prime})f_{B^{\prime}}(B,B^{\prime})]\\ &+\mathbb{E}_{AA^{\prime}}[f_{A}(A,A^{\prime})]\mathbb{E}_{BB^{\prime}}[f_{B}(B,B^{\prime})]\\ &-2\mathbb{E}_{AB}\big{[}\mathbb{E}_{A^{\prime}}[f_{A}(A,A^{\prime})]\mathbb{E}_{Y^{\prime}}[f_{B}(B,B^{\prime})]\big{]},\end{split}$ (10)
where $(\cdot)^{\prime}$ denotes independent copies of variables, and $f_{(\cdot)}$ indicates kernel functions. Efficient calculation of HSIC is often undertaken using an empirical estimate [28]:
$\operatorname{HSIC}(A,B)=\frac{1}{(|A|-1)^{2}}\operatorname{tr}(F_{A}HF_{B}H),$ (11)
where ||$\cdot$|| ⋅ | returns the cardinality. In addition, $H=\mathbf{I}-\frac{1}{|\theta|}\mathbf{11}^{\top}$ is the centring matrix, and $F_{A}=f_{A}(a_{i},a_{j})$ and $F_{B}=f_{B}(b_{i},b_{j})$ are kernel matrices obtained from $a_{i}\in A$ and $b_{i}\in B$.
Input: Blockchain $B$, set of miner nodes $C^{*}$, aggregated model $\bar{\theta}^{t+1}$
Output: Validation result: approved or rejected
1 Retrieve $\bar{\theta}^{t}$ from $b^{t}\in B$;
2 Retrieve set of model parameters $\{\theta_{i}^{t}\}_{i=1}^{r}$ from $B$
3 Calculate $\theta^{t}_{m}\leftarrow\operatorname{Median}(\{\theta_{i}^{t}\}_{i=1}^{r})$
4 Retrieve $h$ from previous $q$ blocks in $B$
5 Estimate $\tau^{t}$ using (13)
6 Calculate $\operatorname{HSIC}(\theta^{t}_{m},\bar{\theta}^{t+1})$;
7 if $\operatorname{HSIC}(\theta^{t}_{m},\bar{\theta}^{t+1})\geq\tau^{t}$ then
8       Approve $\bar{\theta}^{t+1}$
9       Trigger $SC_{3}$ to include HSIC in the next block
10
11 else
12       Reject $\bar{\theta}^{t+1}$
13       Forbid $c_{a}$ and reset the selection process
14       Trigger $SC_{1}$ to select a new aggregator
15
16 end if
Algorithm 2 Post-Aggregation Auditing
In order to use HSIC for measuring statistical independence of $\bar{\theta}^{t+1}$ with available $\theta_{i}^{t}$, we need to minimize the cost of this process, as calculating pairwise HSIC may become computationally burdensome for large values of $r$. Moreover, a threshold needs to be defined to determine whether the estimated HSIC is abnormal. Thus, we choose $\theta^{t}_{m}=\operatorname{Median}(\{\theta_{i}^{t}\}_{i=1}^{r})$ as a representative of all $\theta_{i}^{t}$ and calculate $\operatorname{HSIC}(\theta_{m}^{t},\bar{\theta^{t+1}})$ instead. Another alternative to using the median is calculating the average of $\{\theta_{i}^{t}\}_{i=1}^{r}$. However, the median is more robust against poisonous updates. At each round $t$, a threshold $\tau$ is obtained as:
$h=\bigcup_{t=1}^{q}\operatorname{HSIC}(\theta_{m}^{t},\bar{\theta}^{t+1}),$ (12)
$\tau^{t}=\min(h)-\lambda\sigma,$ (13)
where $h$ is the set of previous $q$ HSIC values, $\lambda$ is the scaling factor, and $\sigma$ is the standard deviation. $\bar{\theta}^{t+1}$ will be approved if $\operatorname{HSIC}(\theta_{m}^{t},\bar{\theta}^{t+1})\geq\tau^{t}$. Note that in (12), only $\operatorname{HSIC}(\theta_{m}^{t},\bar{\theta}^{t+1})$ is estimated in time $t$, as the HSIC for previous time steps is already available in previous blocks.
V-C Blockchain Components
In this work, we implement a simple permissioned blockchain in which participants can join using a private key. An authentication server is considered to facilitate this task. Participants of this network are either users (i.e., formulated as $c_{i}\in C$) or minters (i.e., indicated as $c^{*}_{i}\in C^{*}$). Fig. 1 illustrates the design of TrustChain and shows how blockchain components are orchestrated along with the proposed algorithms.
Blocks
Blocks $b^{t}\in B$ of the DFL contain commonly used header information such as hash to the previous block, timestamp $t$, and block number (i.e., indicated as in the subscript of $b$). Additionally, the header includes a flag that indicates whether this block contains $\theta_{i}^{t}$ or $\bar{\theta}^{t}$. In addition, a signature is included and used for authentication. Moreover, blocks contain data segments consists of $\bar{\theta}^{t}$, $\operatorname{HSIC}(\theta_{m}^{t-1},\bar{\theta}^{t})$, and $\{S_{i}^{t}\}_{i=1}^{q}$. Furthermore, blocks contain SCs that will be triggered when certain conditions are satisfied.
Smart contracts
TrustChain uses three SCs that are triggered when these conditions are met:
  1. 1.
    $SC_{1}$: Miners $c_{i}^{*}\in C^{*}$ request issuing $\bar{\theta}^{t+1}$ for $b^{t+1}$. This triggers PSE (Algorithm 1) to select $c_{a}$.
  2. 2.
    $SC_{2}$: $c_{a}$ requests pushing $\bar{\theta}^{t+1}$ to $b^{t+1}$, which triggers PAA (Algorithm 2) to verify the issued update before adding it to $b^{t+1}$. If the condition is not met, $c_{i}$ corresponding to $c_{a}$ will not be selected for several rounds, and $SC_{1}$ is repeated.
  3. 3.
    $SC_{3}$: If $\bar{\theta^{t+1}}$ is approved, the estimated $\operatorname{HSIC}(\theta_{m}^{t-1},\bar{\theta}^{t})$ in $SC_{2}$ will be included in $b^{t+1}$.
VI Experimental Results
This section first explains the experimental setup used in our experiments and then evaluates the proposed DFL under several scenarios. The experiments are conducted in three phases. First, we evaluate the overall accuracy when TrustChain is employed. Then, the sensitivity to attacks is discussed. Finally, we study PSE and PAA by isolating each component while using TrustChain.
VI-A Experimental Setting
Datasets
Experiments are evaluated using commonly used benchmark datasets in this domain, such as MNIST [29], Fashion-MNIST [30], and CIFAR10 [31] datasets. To create federated datasets, datasets are shuffled at first and then sub-datasets are created by random sampling without replacement. The resulting data is highly imbalanced, and the created subsets have an equal number of samples.
Local models
The architecture of local models differs for each dataset. For experiments on MNIST and Fashion-MNIST, a Multi-Layer Perceptron (MLP) with two hidden layers is used. For CIFAR10, on the other hand, we use a Convolutional Neural Network (CNN) with six convolutional layers followed by two dense layers. Convolution layers are paired with batch normalization, and after every two convolutional layers, max pooling and dropout (ratio=$25\%$) are used. In both networks, hidden layers use ReLU, and the final layer is activated using Softmax. Model parameters are optimized using stochastic gradient descent. The batch size is set to 64 for MLP and 32 for CNN.
DFL parameters
The DFL undergoes 1000 training iterations ($1\leq t\leq 1000$). The number of participants in $C$ is set to $r=100$. We assume $C^{*}\subset C$ and set $k=20$. The window length is set to $q=15$. The scaling factor in (13) is empirically selected and set to $\lambda=0.5$. To facilitate the estimation of $\tau$, we assume the first 50 DFL iterations are safe, and attacks begin at $t=51$.
Attacks scenarios
Using label flipping, the local data on the attacker’s side is poisoned with $\gamma=1$. Local training on the poisoned data results in malicious parameters $\tilde{\theta}$ which are sent to the DFL. Sign flipping is performed by training on clean local data and reversing the sign of gradients. GM is implemented by adding Gaussian noise to $\theta_{i}$ with a mean and standard deviation equal to $0.1$. For TOM, we set $\zeta=0.1$ in the corrupted loss function.
Evaluation metrics
The overall performance of the DFL structure is evaluated in terms of accuracy. Experiments are repeated ten times and the results are averaged. To evaluate PSE and PAA, we measure precision, as it takes both the success rate (true positives) and false alarms (false negatives) into account.
Computing infrastructure
Results were obtained on a computer equipped with an NVIDIA RTX 3080 GPU, an Intel Core i7-12700 CPU, and 32 GB of RAM. The experiments were conducted using Python on an Ubuntu kernel, accessed via the Windows Subsystem for Linux.
VI-B Results Analysis
Figure 2: Classification accuracy of DFL under three different conditions: (a–c) aggregator is benign, (d–f) uses TrustChain and aggregator is not trusted, (g–i) TrustChain is not used and aggregator is not trusted. For each scenario, experiments are separated based on the utilized datasets.
Overall performance
In order to determine the effectiveness of TrustChain, we first compare the classification accuracy of DFL under four different attacks. In these experiments, malicious users attempt to inject poisonous updates into the DFL network. When DFL is not paired with TrustChain, aggregator nodes are selected randomly. Since the DFL is equipped with a robust aggregator, the effect of these attacks is expected to be partially mitigated when not initiated by the aggregator. Here, we select TrimmedMean [32] for the task of robust aggregation. Nevertheless, when these nodes are selected as the aggregator, they will have the opportunity to bypass the robust aggregator. Fig. 2 compares the classification accuracy of TrimmedMean under three conditions when poisoning attacks are launched with different ratios of Byzantine nodes, ranging from five to twenty percent of the nodes’ population. To identify a baseline, we first investigate a scenario where the aggregator is benign and TrustChain is not used in the DFL, as shown in panels (a–c) of Fig. 2. This shows the performance of TrimmedMean in attack scenarios, without a malicious aggregator becoming involved. As expected, the accuracy deteriorates as the ratio of Byzantine nodes increases. Next, we assume the aggregator is not trusted, that is a malicious node can be chosen to carry out the aggregation. Fig. 2(d–f) shows the evaluation results of TrustChain under this condition. The accuracy obtained in these panels is slightly lower than that of Fig. 2(a–c). This indicates that TrustChain can robustly mitigate malicious aggregators, albeit not completely. To reveal the effectiveness of TrustChain in this scenario, we remove it from the experiments in Fig. 2(g–i) to demonstrate the potency of malicious aggregators when they are not audited. As illustrated, a noticeable performance drop has taken place as the bypassed poisoned update rapidly propagates through the DFL network. This effect becomes more pronounced as the ratio of Byzantine nodes increases because malicious aggregators will have a higher chance of being selected in each round.
Sensitivity to attacks
From Fig. 2(g–i), it can be inferred that the portion of Byzantine nodes has a greater impact on their potency compared to the case when TrustChain is used. However, this observation is less noticeable for label flipping, as it varies within a smaller range. Moreover, TOM appears to be more difficult to mitigate, as it results in the lowest classification accuracy in most experiments shown in Fig. 2. For large numbers of Byzantine nodes, sign flipping can also lead to catastrophic errors, when TrustChain is not employed. Nevertheless, the results indicate that TrustChain can mitigate the effect of the Byzantine population to a large extent. This can be seen by comparing panels (d–f) and (g–i) in Fig. 2.
Abblation study
Figure 3: Effect of $q$ on the precision of PSE and PAA. Each component is studied separately, and the precision is averaged for all attack scenarios and datasets.
To study the efficacy of each component, we conduct two sets of experiments where only the selected component is exposed to adversaries. To test the PSE component, we launch the selected attacks from the participants’ side and monitor the nodes PSE selects in this process. We estimate the precision by considering the selection of benign nodes as true positives and malicious nodes as false positives. Similarly, we use the same metric to evaluate PAA by considering correct and false verification as true positives and false positives, respectively. In this process, $\bar{\theta}$ is only corrupted by the aggregator, and nodes do not exhibit malicious behavior before being selected. Fig. 3 illustrates the precision recorded for PSE and PAA. The reported precision is averaged over the results of all attacks and datasets to capture the overall performance of these components. Based on this figure, it can be concluded that a small size of $q$ may lead to unreliable performance for both methods. For PSE, larger values of $q$ result in better precision because analyzing the drift over longer intervals is more accurate. However, the precision does not significantly change after a certain point. On the other hand, we notice that PAA is more sensitive to the choice of $q$. The plotted curve indicates that PAA works best for an optimal value of $q$, and values that are too small or too large result in deteriorated precision. This is due to the fact that when the window length is too small, PAA struggles to find an appropriate value for $\tau$. On the other hand, when $q$ is too large, older HSIC values that are too old are not removed from the window, leading to an inaccurate estimation of $\tau$, which only decreases and does not increase.
Significance test
Fig. 4 shows Critical Difference (CD) diagrams obtained from the post-hoc Friedman test. The significance level (i.e., parameter $\alpha$) is set to $0.05$ in this test. This test estimates the significance of differences among the results obtained from each method across all experiments. Based on the determined CD level of this test, methods that are not significantly different in terms of accuracy are connected and grouped using colored lines. The illustrated CD diagram shows that employing TrustChain results in a performance that is statistically similar to that having a benign aggregator. Moreover, the proposed method significantly boosts the performance when the aggregator is not trusted (indicated as DFL in Fig. 4).
Figure 4: Critical difference diagram obtained from the post-hoc Friedman test. The significance level is set to $0.05$.
VII Conclusion
This paper addressed the issue of malicious aggregators in DFL systems by proposing a blockchain-enabled solution for trustworthy aggregation, called TrustChain. The proposed method eliminates malicious aggregators in two steps. Firstly, the miners participating in the aggregation process are scored based on their tendency to drift from the distribution of aggregated parameters in previous DFL iterations. This step excludes miners who exhibit suspicious behavior. Secondly, the aggregated updates are checked before being pushed to the DFL by measuring statistical independence using HSIC between the aggregated model and the updates from this round. Aggregated models are approved based on a threshold that is dynamically adjusted according to the information available in previous blocks of the blockchain. This step adds an extra layer of security against trusted nodes that become rogue upon being selected as the aggregator. The proposed method was evaluated on three benchmark datasets and against four poisoning attacks, with varying numbers of Byzantine nodes. The results demonstrated the robustness of TrustChain in eliminating malicious aggregators. Future research directions will focus on studying TrustChain under more sophisticated attacks and integrating it with privacy-preserving methods to extend its applicability.